the first side of a triangle is 3cm longer than the second side. the third side is twice the second side. if the perimeter is 59cm, what is the length of each side?
Answer by bam878s(77)

Let x = length of the second side, y=length of the first side, and z=length of the third side.
From the question we have,
1.x=length of the second side
2.y=length of the first side = x+3.
3.z=length of the third side = 2x.
4.The sum of these sides, by definition, is the perimeter=59cm
So,
59 = x + (x+3) + (2x).
59 = x+x+3+2x.
59=4x+3.
solve for x.
59-3=4x+3-3
56=4x.
divide by 4 on both sides.
14=x.
Now plug x back into the other formulas.
y = length of the second side = x+3 = 14+3 = 17.
z = length of the thrid side = 2x = 2(14) = 28.